The following frequency polygon displays number of vehicles sold in various price ranges. Use this frequency polygon to answer the questions.

Find the percentage of vehicles sold for between $24999.50$⁢24999.50 and $34999.50$⁢34999.50. Round your answer to the nearest percent.

The following frequency polygon displays number of vehicles sold in various price ranges. Use this frequency polygon to answer the questions. Prices of New Vehicles Sold by a Dealership 90 1 74 75 63 60 45 36 26 30 15 $7499.50 S12,499.50S17,499.50S22,499.50S27,499.50S32,499.50S37,499.50$42,499.50 Prices Step 5 of 5: Find the percentage of vehicles sold for between $24999.50 and $34999.50. Round your answer to the nearest percent. Answer Tables Keypad How to enter your answer Previous Step Answer(s) Submit Answer
Answer #1

total frequency =63+32+26+36+39+74=270

number of vehicles between 24999.5 and 34999.50 =36+39=75

hence % of vehicles between 24999.5 and 34999.50 =(75/270)*100 =27.78% =28%